11x-6-4x+1=9x-8-2x+12
Answer by soccerdude716(1)

11x-6-4x+1=9x-8-2x+12
This one is kinda tricky so you have to pay attention to what you do. First you are going to combine like terms on both sides so this is what you get:
11x-6-4x+1
you combine the x's. it is easier to solve if you add the opposites, so you get
11x+(-6)+(-4x)+1
11x+-4x=7x
Put it back into the equation
7x+(-6)+1
Now you want to combine the numbers -6 and 1 and get -5
so you first combined equation is 7x+(-5)
Now you do the same thing, combine variables than rest of equation.
9x-8-2x+12 add the oppostie 9x+(-8)+(-2x)+12
variables= 9x+(-2x)= 7x
rest= (-8)+ 12 = 4
so you get 7x+4
put it together and you got 7x+(-5)=7x+4

now you are going to solve for x. First you add the -5 and 4 and get -1 on one side. and subtract 7x and put it on the other side and get 0x=-1. This is impossible so your answer is no solution